An angles measure is 6 degrees more than 3 times the measure of its complement. Find th measure of the angle.
Answer by checkley79(3341)

x+y=180
x=3y+6
(3y+6)+y=180
47=180-6
4y=174
y=174/4
y=43,5 degrees for the smaller angle.
x+43.5=180
x=180-43.5
x=136.5 ans. for the largest angle.
Proof:
43.5+136.5=180
180=180